OpenStudy (anonymous):

Use the substitution method to solve x^2+2y^2=9 x-y+3=0 can someone show me the steps?

OpenStudy (anonymous):

I have already solved this before. You've asked the same question lol, search for the old one ^_^

OpenStudy (anonymous):

oh wait, it's different ^^" my bad.

OpenStudy (anonymous):

yea the y^2 is throwing me off.

OpenStudy (anonymous):

i came up w/ x=(-1,-3) y=(2,0) but im not sure thats right...

OpenStudy (anonymous):

Alright, first of all, call your equations as (a) and (b). Now, let's re-write equation (b) in terms of y and we'll get: y = 3 + x Substitute this equation in (a) and you'll get: \[x^2 + 2(3+x)^2 - 9 = 0\] \[x^2 + 2(3+x)(3+x) -9 =0\] \[x^2 + 2x^2 + 12x + 18 -9 = 0\] \[3x^2 + 12x -9 = 0\] \[(3x-3)(x-9) = 0\] so x = 1 and x =3 Plug these 2 x's in equation b and find the 2 points ^_^ give it a try now :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Your points should be (1,4) and (3, 6) :)

OpenStudy (anonymous):

DOn't you mean x = 1 and x = 9

OpenStudy (anonymous):

lol right right! >_< thank you :)

OpenStudy (anonymous):

wait no!

OpenStudy (anonymous):

oh dear , I messed up again the equation should be: (3x-3)(x-3) = 0 instead of (x-9)

OpenStudy (anonymous):

so you have x = 1 and x = 3 ^^" sorry lol

OpenStudy (anonymous):

plug these in equation (b) and you'll get (1,4) and (3,6) as your points :)

OpenStudy (anonymous):

and thank you blexting :)

OpenStudy (anonymous):

it wouldnt be (1,4) and (3,12)?

OpenStudy (anonymous):

no ,it'll be (1,4 ) and (3,6) I made a silly mistake in the previous last line where I have put: (3x -3)(x-9) = 0 it should be : (3x -3)(x-3) = 0 instead :)

OpenStudy (anonymous):

plug x = 3 in the equation and you'll get the other point (3,6) ^_^

OpenStudy (anonymous):

ok I see. that x-9 was throwing me off. Thans for your help!

OpenStudy (anonymous):

np :) again, sorry for the mess ^^"

OpenStudy (anonymous):

Wait. x^2+2x+12+18-9=0 wouldnt this simplify to 3x^2+2x+9=0 ?

OpenStudy (anonymous):

you have negative 9

OpenStudy (anonymous):

oh right! >_< again sorry lol

OpenStudy (anonymous):

yo'll have negative 3 and not 9

OpenStudy (anonymous):

Ok I think i have it.

OpenStudy (anonymous):

awesome! :)

OpenStudy (anonymous):

awesome indeed

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

when i solve I keep coming up with x=(-1,-3) y=(2,0)?

OpenStudy (anonymous):

y=2 and x=1

OpenStudy (anonymous):

or y=0 and x=3

OpenStudy (anonymous):

yes! excellent js! :)

OpenStudy (anonymous):

I'll give you a star ^_^

OpenStudy (anonymous):

truly excellent...

OpenStudy (anonymous):

a star from me too

OpenStudy (anonymous):

lol ^_^

OpenStudy (anonymous):

i mean y=0 and x=-3

OpenStudy (anonymous):

lol abigail :)

OpenStudy (anonymous):

i know what you mean, dont wory

OpenStudy (anonymous):

thank you for your support too ^_^

OpenStudy (anonymous):

you know what I mean?

OpenStudy (anonymous):

i know what i know

OpenStudy (anonymous):

so, what did I mean?

OpenStudy (anonymous):

I know what you know I know

OpenStudy (anonymous):

._. ...now I don't even know what I have said

OpenStudy (anonymous):

i know you dont know what you think you dont know

OpenStudy (anonymous):

I don't even know what I know now x_x

OpenStudy (anonymous):

LOL, stop my head is spinning! @_@

OpenStudy (anonymous):

Well guys i know i appreciate the help:)

OpenStudy (anonymous):

np js :)

OpenStudy (anonymous):

i know it is

OpenStudy (anonymous):

I don't even know what is what right now, but I do know that the word "know" sounds funny nowx_x

OpenStudy (anonymous):

just say "know" 29 times out loud, and then pause for 5 sec, and say it again, and i promisse, it will sound weird

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

go finish your 5 $ work =P

OpenStudy (anonymous):

finished all the orders already

OpenStudy (anonymous):

made a fortune

OpenStudy (anonymous):

atleast you had fun :)

OpenStudy (radar):

This solution was very difficult to follow. I hope js14 understands the final answer is x=-3, y=0